OpenStudy (anonymous):

Choose two axioms that allow 6 + (x + 5) to be written x + 11. commutative - addition distributive associative - multiplication symmetric commutative - multiplication associative - addition identity - addition

OpenStudy (zzr0ck3r):

6+(x+5) = 6+(5+x) because of commutability 6+(5+x) = (6+5) + x = 11+x because of associativity.

OpenStudy (anonymous):

Thanks but I had it figured out. Still thank you